OpenStudy (anonymous):

if the product of 10 times 10 times 10 is a 1000 then what three prime numbers whose product is 1001

OpenStudy (anonymous):

@mathstudent55 Help me please?

OpenStudy (whpalmer4):

Well, if 10*10*10 = 1000, can you think of 3 prime numbers each in the ballpark of 10?

OpenStudy (whpalmer4):

Do you know what a prime number is?

OpenStudy (anonymous):

A number that cannot be divided by any other.

OpenStudy (whpalmer4):

So, can you list the first 10 or so?

OpenStudy (anonymous):

Me, it's his question.

OpenStudy (whpalmer4):

Anyone? Bueller? Bueller? :-)

OpenStudy (mathstudent55):

1001 is not divisible by 3 Since there is no divisibility test for 7, try dividing by 7

OpenStudy (anonymous):

factor 3 and 5 don't work, check 7

OpenStudy (anonymous):

actually there is an easy trick to see that this is divisible by 11 start with the last digit, then compute as follows: \[1-0+0-1=0\] so the remainder when dividing by 11 is 0, i.e. it is divisible by 11

OpenStudy (anonymous):

11, 7, 13

OpenStudy (anonymous):

YES !!!!!!!!!!!!!!!1

OpenStudy (anonymous):

11 * 7 * 13 = 1001

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

Can I have a medal?

OpenStudy (anonymous):

i dont know how to give them

OpenStudy (anonymous):

Click the blue best response button to the right of my name

OpenStudy (anonymous):

k

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

thx

OpenStudy (whpalmer4):

Alternatively, look at the list of the first 10 primes: 2, 3, 5, 7, 11, 13, 17, 19, 23, 29 Pick the three closest to 10, multiply them together and make sure the product is 1001.